# Dynamic Programming for Symbolic Boolean Realizability and Synthesis

### v1

@@ 1 Introduction
Motivated by applications that require the computation of both realizability sets and witness functions, we describe here a decision-diagram approach that can also handle partially realizable specifications, where the realizability set is neither empty nor necessarily universal. (Indeed, in our benchmark suit, about 30% of the benchmarks are partially realizable.) Our tool, DPSynth computes the realizability sets and witness functions with respects to these sets. In contrast, prior tools, such as [1, 17, 18], focus on fully realizable specifications, where the realizability sets are universal, and return "Unrealizable" for partially realizable specification. This reduced scope of applicability enables these tools to be faster.
@@ 5 Experimental Evaluation
Our first experiment verifies whether DPSynth is competitive with other non-BDD-based synthesis solvers, as motivated in Section 1. For a comparison with Manthan [17], a leading tool that is not based on decision diagrams, DPSynth has superior performance.
We present a general picture by the following table, which shows the overall strength of the dynamic-programming decision-diagram approach, by measuring the time and space usage of experiments on the dataset selected from QBFEVAL’16 to QBFEVAL’19. DPSynth and Manthan [17] each solves for some benchmarks that the other does not solve.
In order to compare the running time, we compare against Manthan using the DPSynth end-to-end synthesis time, which is the sum of tree-decomposition time, compilation, realizability-checking time, and synthesis time. For partially realizable instances, since Manthan returns UNSAT for partially realizable benchmarks, we compare with Manthan by measuring the realizability time of DPSynth, which is the sum of tree-decomposition, compilation and realizability-checking time.
As the overall picture of synthesis solving, DPSynth shows a better time performance on almost all instances. DPSynth and Manthan each has strength on some instances, but the number of benchmarks that only DPSynth solves is larger than those solved only by Manthan. On the intersection of solved sets, DPSynth takes less time usage. See Appendix 0.D for more illustrative data on specific fully and partially realizable benchmarks.
Number of benchmarks
solved by Manthan 79
solved by DPSynth 102
solved by both Manthan and DPSynth 70
solved by Manthan but not by DPSynth 9
solved by DPSynth but not by Manthan 32
Next, the rest of this section focuses on comparative performance of this tool against another BDD-based approach.
@@ 5.1 Realizability-Checking Phase
Methodology: To examine our dynamic-programming graded-project-join-tree-based approach for boolean realizability, we developed a software tool, DPSynth, which implements the theoretical framework described above. As explained in the Introduction, it makes sense to compare DPSynth to Factored RSynth, rather than to tools that cannot handle partially realizable specifications. Prior work [27] already demonstrated that RSynth practically never outperforms Factored RSynth, while Factored RSynth typically outperforms RSynth, so we compare in this paper to Factored RSynth. (See Appendix D for a comparison to a non-decision-diagram tool.) We measured the time and space performance for determining realizability on a set of mature benchmarks, described below. The experiments aim at answering the following research questions:
@@
Figure 2(a) shows overall running-time comparison between DPSynth and Factored RSynth for realizability checking. A clear pattern that emerges is there is a difference in relative performance between very small problems (solvable within $1$ millisecond) and larger problems. DPSynth underperforms Factored RSynth on small instances, but outperforms on larger instances and the difference increases exponentially as input size grows. We conclude that planning overhead dominates on small input instances, but that effect fades off as instances get larger and the planning pays off. In memory usage comparison [^3] The charts for space consumption for both phases are in the Appendix. 0.C. – using peak node count as a measurement – DPSynth uses less memory than Factored RSynth. Here the graded project-join-tree approach is advantageous, and planning incurs no overhead.
@@ Appendix 0.C Space Performance Charts
Benchmark Manthan time (s) DPSynth time (s) Manthan / DPSynth time
@@
Benchmark Manthan time (s) DPSynth time (s) Manthan / DPSynth time
@@
factorization8 26.02147222 0.00651449 3994.398981
rankfunc37_signed_16 2.976966858 0.009265158 321.3077271
stmt137_903_911 18.31892133 0.0820988 223.132632
stmt19_3_214 67.52295566 410.1905171 0.164613644
stmt19_66_214 84.54595137 41.93132509 2.016295721
stmt2_976_999 3228.63344 13.59846008 237.4264014
stmt21_5_134 47.65692377 3.044376174 15.65408512
stmt21_84_215 88.41444421 37.0518531 2.386235419
stmt27_16_97 53.51791215 1.509911213 35.44441003
stmt27_584_603 24.56984329 0.119144383 206.2190653

### v2

# Yi Lin 11 $0$-0001-8443-2246    Lucas M. Tabajara Currently at Runtime Verification, Inc.11 $0$-0001-9608-1404    Moshe Y. Vardi 11 $0$-0002-0661-5773

@@ 1 Introduction
Motivated by applications that require the computation of both realizability sets and witness functions, we describe here a decision-diagram approach that can also handle partially realizable specifications, where the realizability set is neither empty nor necessarily universal. (Indeed, in our benchmark suit, about 30% of the benchmarks are partially realizable.) Our tool, DPSynth computes the realizability sets and witness functions with respects to these sets. While several recent synthesis tools do provide witnesses for partially-realizable specifications, cf. [1, 17, 18], not all of them directly output the realizability set, requiring it, instead, to be computed from the witnesses and the original formula.
@@ 5.1 Realizability-Checking Phase
Methodology: To examine our dynamic-programming graded-project-join-tree-based approach for boolean realizability, we developed a software tool, DPSynth, which implements the theoretical framework described above. We choose to compare DPSynth to Factored RSynth, which as explained in the introduction is the closest existing tool, also being based on decision diagrams and outputting the realizability set along with the witnesses. Prior work [27] already demonstrated that RSynth practically never outperforms Factored RSynth, while Factored RSynth typically outperforms RSynth, so we compare in this paper to Factored RSynth. (See Section 5.4 for an additional comparison to a non-decision-diagram tool.) We measured the time and space performance for determining realizability on a set of mature benchmarks, described below. The experiments aim at answering the following research questions:
@@
Figure 2(a) shows overall running-time comparison between DPSynth and Factored RSynth for realizability checking. A clear pattern that emerges is there is a difference in relative performance between very small problems (solvable within $1$ millisecond) and larger problems. DPSynth underperforms Factored RSynth on small instances, but outperforms on larger instances and the difference increases exponentially as input size grows. We conclude that planning overhead dominates on small input instances, but that effect fades off as instances get larger and the planning pays off. In memory usage comparison [^4] The charts for space consumption for both phases are in the Appendix C. – using peak node count as a measurement – DPSynth uses less memory than Factored RSynth. Here the graded project-join-tree approach is advantageous, and planning incurs no overhead.
@@ 5.3 Tree Widths and Realizability
5.4 Comparison with Non-BDD-Based Synthesis
To complement out evaluation, we include additional experiments to verify whether DPSynth is competitive with non-BDD-based synthesis solvers, as motivated in Section 1. We compare with Manthan [17], a leading tool that is not based on decision diagrams, and find that DPSynth performs favorably.
We present a general picture by the following table, which shows the overall strength of the dynamic-programming decision-diagram approach, by measuring the time and space usage of experiments on the dataset selected from QBFEVAL’16 to QBFEVAL’19. DPSynth and Manthan each is able to solve some benchmarks that the other does not solve.
In order to compare the running time, we compare against Manthan using the DPSynth end-to-end synthesis time, which is the sum of tree-decomposition time, compilation, realizability-checking time, and synthesis time.
As the overall picture of synthesis solving, DPSynth shows a better time performance on most instances. DPSynth and Manthan each has strength on some instances, but the number of benchmarks that only DPSynth solves is larger than those solved only by Manthan. On those that are solved by both, DPSynth takes less time in most of them. See Appendix D for more illustrative data on specific fully and partially-realizable benchmarks.
Number of benchmarks
solved by Manthan 79
solved by DPSynth 102
solved by both Manthan and DPSynth 70
solved by Manthan but not by DPSynth 9
solved by DPSynth but not by Manthan 32
@@ Appendix 0.C Space Performance Charts
Benchmark Manthan time (s) DPSynth time (s) Manthan / DPSynth time
@@
factorization8 26.02147222 0.01355423 1919.804534
rankfunc37_signed_16 2.976966858 0.020192158 147.4318326
stmt137_903_911 18.31892133 0.2235378 81.94999382
stmt19_3_214 67.52295566 970.6645171 0.069563639
stmt19_66_214 84.54595137 148.8763251 0.567893863
stmt2_976_999 3228.63344 57.85746008 55.80323498
stmt21_5_134 47.65692377 14.28987617 3.335013067
stmt21_84_215 88.41444421 135.9611531 0.650291956
stmt27_16_97 53.51791215 9.510231213 5.627403893
stmt27_584_603 24.56984329 0.329176383 74.64035867